1) . A passenger on a moving train walks at a speed of 1.90 m/s due north relative to the train. The passenger's speed with respect to the ground is 4.5 m/s at an angle of 31.0° west of north. What are the magnitude and direction of the velocity of the train relative to the ground?

2) A ship sets sail from Rotterdam, The Netherlands, heading due north at 8.00 m/s relative to the water. The local ocean current is 1.55 m/s in a direction 40° north of east. What is the velocity of the ship relative to the earth?
____ m/s  

____° N of E

3)

A woman starts to swim directly across a 8.0-km-wide river. Her speed with respect to the water is 2.3 m/s. The river current carries the woman downstream at a speed of 0.91 m/s.

(a)

How much time, in seconds, does it take her to cross the river?

(b) How far downstream, in meters, will the river carry her by the time she reaches the other side of the river?

_____ meters
Answer #1

1)

VPG = VPT + VRG

VPG = 4.5i + 1.9j

|VPG| = sqrt(4.5^2 + 1.9^2) = 4.88 m/s

Direction, theta = arctan(1.9/4.5) = 22.89 degree North of East

Comment in case any doubt please rate my answer....

According to the HOMEWORKLIB RULES I can answer one question at a time so please post other questions seperately .....